Give the oxidation state of all the atoms in the following: HPO4- H2O2 ClO4-

  1. Give the oxidation state of all the atoms in the following:
    1. HPO4-
    1. H2O2
    1. ClO4-
Answer #1

a. Oxidation state of H =+1

Oxidation state of O = - 2

Oxidation state of P =x

Now,

1 + x + 4(-2) = -1

1+x-8=-1

x= -1+7

x=+6

Thus, oxidation state of P =6

b. Oxidation state of H =+1

Oxidation state of O =x

Now,

2(1) +2(x) =0

2+2x=0

x= -2/2

x=-1

Thus, oxidation state of O =-1

c. Oxidation state of O =-2

Oxidation state of Cl =x

Now,

x+4(-2) =-1

x-8=-1

x=-1+8

x=7

Thus, Oxidation state of Cl= 7